# Upper bound for the total mean curvature of spin fill-ins

### v1

@@ Abstract.
Gromov conjectured that the total mean curvature of the boundary of a compact Riemannian manifold can be estimated from above by a constant depending only on the boundary metric and on a lower bound for the scalar curvature of the fill-in. We prove Gromov’s conjecture if the manifolds are spin and the mean curvature is non-negative.
@@ 1. Introduction
for spin manifolds. This follows by combining an upper Dirac eigenvalue estimate by the author ([Baer]*Main Theorem) with a lower one by Hijazi, Montiel, and Roldán ([HMR]) as has been observed by Brendle, Tsiamis, and Wang in [BTW]. See also [CHZ]*Theorem 1.5 by Cecchini, Hirsch, and Zeidler for (1) in the case $\sigma=0$. Clearly, estimate (1) is only meaningful if the mean curvature is positive.
@@
Let $X$ be a compact Riemannian spin manifold of dimension $n+1\geq 2$ with smooth boundary $\partial X=M$. Let $\lambda\geq 0$ be such that the scalar curvature of $X$ satisfies $\mathrm{scal}_{X}\geq-\lambda^{2}$. Assume that the mean curvature $H$ of the boundary satisfies $H\geq 0$. Then we have
$\int_{M}H\leq C(M)+\sqrt{\tfrac{n}{n+1}}\,\lambda\,\mathsf{vol}(M),$
@@
If $\lambda=0$, $H>0$, $X$ is spin, and each connected component of $M$ admits an isometric embedding as a strictly convex hypersurface in $\mathbb{R}^{n+1}$, then this has been proved by Shi and Tam in [ST]*Theorems 1 and 4.1. The constant $C(M)$ is in this case given by the total mean curvature of the isometric embedding of $M$ into $\mathbb{R}^{n+1}$. The condition that each connected component of $M$ embeds isometrically as a strictly convex hypersurface into $\mathbb{R}^{n+1}$ has been relaxed by Eichmair, Miao, and Wang in [EMW] to require the components to have non-negative scalar curvature and admit embeddings as star-shaped hypersurfaces.
@@
If $M=T^{n}$ is a torus, $X=D^{2}\times T^{n-1}$, $n\leq 6$, and $H>0$, then Gromov’s conjecture has been proved by Wang in [W]*Theorems 1.6 and 1.8. If the torus $M$ carries a flat metric, then $C(M)$ is related to the systole of $M$.
Bernhard Hanke has informed me that, if the mean curvature is strictly positive, surgery and deformation methods can be used to derive an upper bound on the total mean curvature that depends on $M$, $\lambda$, and a positive lower bound on $H$ ([H]).
@@ 2. A boundary value problem
are Fredholm. If $\Phi\in H^{1}(X,\Sigma X)$ lies in the kernel of $(D-i\lambda)\oplus P^{+}(\cdot|_{M})$, then $(D-i\lambda)\Phi=0$ and $P^{+}(\Phi|_{M})=0$. By Lemma 1, we then also have $P^{-}(\Phi|_{M})=0$. Hence, $\Phi|_{M}=0$. By the unique continuation property for Dirac-type operators, we conclude $\Phi=0$ (see e.g. the proof of [BBHW]*Corollary B.2). Thus, the operator
@@ 4.2. The estimate
$0$ $=\int_{X}\Big\langle\Big(D-i\sqrt{\tfrac{n+1}{n}}\tfrac{\lambda}{2}\Big)\Phi_{j},\Big(D-i\sqrt{\tfrac{n+1}{n}}\tfrac{\lambda}{2}\Big)\Phi_{j}\Big\rangle$
$=\int_{X}\Big\langle\Big(D+i\sqrt{\tfrac{n+1}{n}}\tfrac{\lambda}{2}\Big)\Big(D-i\sqrt{\tfrac{n+1}{n}}\tfrac{\lambda}{2}\Big)\Phi_{j},\Phi_{j}\Big\rangle$
@@
$\leq\frac{1}{2}\sum_{j=1}^{m}\int_{M}H|\Phi_{j}|^{2}$ (6)
$\leq\sum_{j=1}^{m}\Big[\int_{M}\langle D_{M}\,\Phi_{j},\Phi_{j}\rangle+\sqrt{\frac{n}{n+1}}\frac{\lambda}{2}\int_{M}|P^{+}\Phi_{j}|^{2}\Big].$ (7)
In (6) we used that $H\geq 0$. For the first term we find
@@
where we applied Lemma 1 in (8). For the second term we get
@@
Combining (7), (9), and (10) we obtain
$\int_{M}H\leq C(M)+\sqrt{\frac{n}{n+1}}\lambda\mathsf{vol}(M).$
@@ 5. Concluding remarks
The condition that the mean curvature is non-negative has been used only in inequality (6). It is unclear to the author whether this condition can be dropped. One would expect that $H$ being negative somewhere on $M$ would make $\int_{M}H$ smaller and thus easier to estimate from above. Note that the Gauss-Bonnet argument for $\dim(M)=1$ in the introduction does not require any positivity assumption on $H$.
@@
The definition of the constant $C(M)$ in Section 4.1 depends on the spin structure of $M$. However, $M$ being compact, it has only finitely many spin structures. Thus, $C(M)$ can be chosen independently of the spin structure.

### v2

@@ Abstract.
Gromov conjectured that the total mean curvature of the boundary of a compact Riemannian manifold can be estimated from above by a constant depending only on the boundary metric and on a lower bound for the scalar curvature of the fill-in. We prove Gromov’s conjecture if the manifolds are spin and the mean curvature $H$ is non-negative. One can also allow $H$ to take negative values, but then the constant depends on a (negative) lower bound for $H$.
@@ 1. Introduction
for spin manifolds. This follows from [HMR]*Theorem 2 by Hijazi, Montiel, and Roldán. Here the constant $C(M,\sigma)$ depends on the first non-negative eigenvalue of the Dirac operator of $M$ and on $\sigma$. Combining this with an upper Dirac eigenvalue estimate by the author ([Baer]*Main Theorem), one obtains (1) where the constant depends on the hyperspherical radius of $M$ and on $\sigma$, as has been observed by Brendle, Tsiamis, and Wang in [BTW]. See also [CHZ]*Theorem 1.5 by Cecchini, Hirsch, and Zeidler for (1) in the case $\sigma=0$. Clearly, estimate (1) is only meaningful if the mean curvature is positive.
@@
Let $X$ be a compact Riemannian spin manifold of dimension $n+1\geq 2$ with smooth boundary $\partial X=M$. Let $\lambda\geq 0$ be such that the scalar curvature of $X$ satisfies $\mathrm{scal}_{X}\geq-\lambda^{2}$. Let $\eta\geq 0$ be such that the mean curvature $H$ of the boundary satisfies $H\geq-\eta$. Then we have
$\int_{M}H\leq C(M)+\Big(\sqrt{\tfrac{n}{n+1}}\,\lambda+\eta\Big)\mathsf{vol}(M),$
@@
All previously known results need the assumption that $H\geq 0$ or even $H>0$. If $\lambda=0$, $H>0$, $X$ is spin, and each connected component of $M$ admits an isometric embedding as a strictly convex hypersurface in $\mathbb{R}^{n+1}$, then this has been proved by Shi and Tam in [ST]*Theorems 1 and 4.1. The constant $C(M)$ is in this case given by the total mean curvature of the isometric embedding of $M$ into $\mathbb{R}^{n+1}$. The condition that each connected component of $M$ embeds isometrically as a strictly convex hypersurface into $\mathbb{R}^{n+1}$ has been relaxed by Eichmair, Miao, and Wang in [EMW] to require the components to have non-negative scalar curvature and admit embeddings as star-shaped hypersurfaces.
@@
Mantoulidis and Miao investigated in [MM] the case where $n=2$ and $M$ has arbitraty topology, again if $\lambda=0$ and $H>0$.
If $M=T^{n}$ is a torus, $X=D^{2}\times T^{n-1}$, $n\leq 6$, and $H>0$, then Gromov’s conjecture has been proved by Wang in [W]*Theorems 1.6 and 1.8. In this case, $\lambda$ is allowed to be negative. If the torus $M$ carries a flat metric, then $C(M)$ is related to the systole of $M$.
In the mean time, Frenck, Hanke, and Hirsch have presented a different proof of Gromov’s conjecture if $H\geq 0$, using surgery and deformation methods ([FHH]). Their approach has the advantage that it also works in the non-spin case if $n\leq 6$.
@@ 2. A boundary value problem
are Fredholm. If $\Phi\in H^{1}(X,\Sigma X)$ lies in the kernel of $(D-i\lambda)\oplus P^{+}(\cdot|_{M})$, then $(D-i\lambda)\Phi=0$ and $P^{+}(\Phi|_{M})=0$. By Lemma 1, we then also have $P^{-}(\Phi|_{M})=0$. Hence, $\Phi|_{M}=0$. By the unique continuation property for Dirac-type operators, we conclude $\Phi=0$ (see e.g. the proof of [BBHW]*Corollary B.2) [^1] This also follows from (4) unless $\lambda=0$.. Thus, the operator
@@ 4.2. The estimate
$0$ $=\int_{X}\Big\langle\Big(D+i\sqrt{\tfrac{n+1}{n}}\tfrac{\lambda}{2}\Big)\Big(D-i\sqrt{\tfrac{n+1}{n}}\tfrac{\lambda}{2}\Big)\Phi_{j},\Phi_{j}\Big\rangle$
@@
$=\frac{1}{2}\sum_{j=1}^{m}\int_{M}\Big[H|\Phi_{j}|^{2}-H|P^{-}\Phi_{j}|^{2}\Big]$
$\leq\sum_{j=1}^{m}\Big[\int_{M}\langle D_{M}\,\Phi_{j},\Phi_{j}\rangle+\sqrt{\frac{n}{n+1}}\frac{\lambda}{2}\int_{M}|P^{+}\Phi_{j}|^{2}-\frac{1}{2}\int_{M}H|P^{-}\Phi_{j}|^{2}\Big].$ (7)
For the first term we find
@@
where we applied Lemma 1 in (8). For the second term (7) we get
@@
For the third term in (7) we obtain, again using Lemma 1:
$-\frac{1}{2}\sum_{j=1}^{m}\int_{M}H|P^{-}\Phi_{j}|^{2}$ $\leq\frac{\eta}{2}\sum_{j=1}^{m}\|P^{-}\Phi_{j}\|^{2}\leq\frac{\eta}{2}\sum_{j=1}^{m}\|P^{+}\Phi_{j}\|^{2}=\frac{\eta}{2}\sum_{j=1}^{m}\|P^{+}\varphi_{j}\|^{2}=\frac{\eta}{2}\mathsf{vol}(M).$ (11)
Combining (7), (9), (10), and (11) we find
$\int_{M}H\leq C(M)+\bigg(\sqrt{\frac{n}{n+1}}\lambda+\eta\bigg)\mathsf{vol}(M).$
@@ 5. Concluding remarks
It seems strange that one needs a lower bound on the mean curvature in order to obtain an upper bound on its integral. It is unclear to the author whether this dependence can be removed. Note that the Gauss-Bonnet argument for $\dim(M)=1$ and $\lambda=0$ in the introduction does not need a lower bound on $H$.
@@
The definition of the constant $C(M)$ in Section 4.1 depends on the spin structure of $M$. However $M$, being compact, has only finitely many spin structures. Thus, $C(M)$ can be chosen independently of the spin structure.